# Rational Angle Bisection Problem in Higher Dimensional Spaces and Incenters of Simplices over Fields

### v1

@@ Abstract
In this article, we generalize the following problem, which is called the rational angle bisection problem, to the $n$-dimensional space $k^{n}$ over a subfield $k$ of $\mathbb{R}$: on the coordinate plane, for which rational numbers $a$ and $b$ are the slopes of the angle bisectors between two lines with slopes $a$ and $b$ rational? First, we give a few characterizations of when the angle bisectors between two lines with direction vectors in $k^{n}$ have direction vectors in $k^{n}.$ To find solutions to the problem in the case when $k=\mathbb{Q},$ we also give a formula for the integral solutions of $x_{1}{}^{2}+\dots+x_{n}{}^{2}=dx_{n+1}{}^{2}$, which is a generalization of the negative Pell’s equation $x^{2}-dy^{2}=-1,$ where $d$ is a square-free positive integer. Second, by applying the above characterizations, we give a necessary and sufficient condition for the incenter of a given $n$-simplex with $k$-rational vertices to be $k$-rational. On the coordinate plane, we prove that every triangle with $k$-rational vertices and incenter can be obtained by scaling a triangle with $k$-rational side lengths and area, which is a generalization of a Heronian triangle. We also state certain fundamental properties of a few centers of a given triangle with $k$-rational vertices.
@@ 1 Introduction
Throughout this article, let $n\geq 2$ be an integer, and let $k$ be a subfield of $\mathbb{R}.$ We consider the following problem in $\mathbb{R}^{n},$ which is a generalization of the rational angle bisection problem on $\mathbb{R}^{2}$ (see [3]).
@@
For which $\bm{a},$ $\bm{b}\in k^{n}$ do the angle bisectors between two lines with direction vectors $\bm{a}$ and $\bm{b}$ have direction vectors in $k^{n}$?
@@
For which $\bm{a},$ $\bm{b}\in O_{k}{}^{n}$ do the angle bisectors between two lines with direction vectors $\bm{a}$ and $\bm{b}$ have direction vectors in $O_{k}{}^{n}$?
@@
      One of the angle bisectors between two lines with direction vectors $\bm{a}$ and $\bm{b}$ has direction vector $\bm{c}\in k^{n}.$
@@
      One of the angle bisectors between two hyperplanes with normal vectors $\bm{a}$ and $\bm{b}$ has normal vector $\bm{c}\in k^{n}.$
@@
      One of the angle bisectors between two lines with direction vectors $\bm{a}$ and $\bm{b}$ has direction vector $\bm{c}\in O_{k}{}^{n}$.
@@
      One of the angle bisectors between two hyperplanes with normal vectors $\bm{a}$ and $\bm{b}$ has normal vector $\bm{c}\in O_{k}{}^{n}$.
@@
    If $\bm{a}=(2,1,1)$ and $\bm{b}=(3,6,2),$ then the angle bisectors between two lines with direction vectors $\bm{a}$ and $\bm{b}$ have direction vectors $7\bm{a}\pm 3\bm{b},$ since $|\bm{a}|=3$ and $|\bm{b}|=7$ imply that $7\bm{a}$ and $3\bm{b}$ span a rhomb (see Figure 1).
@@
    If $\bm{a}=(1,1,4)$ and $\bm{b}=(3,4,5),$ then the angle bisectors between two lines with direction vectors $\bm{a}$ and $\bm{b}$ have direction vectors $5\bm{a}\pm 3\bm{b},$ since $|\bm{a}|=3\sqrt{2}$ and $|\bm{b}|=5\sqrt{2}$ imply that $5\bm{a}$ and $3\bm{b}$ span a rhomb.
@@
On $\mathbb{R}^{2},$ given triangle $ABC$ over $k$ with side lengths $BC=a,$ $CA=b,$ and $AB=c,$ inradius $r,$ and exradii $r_{A},$ $r_{B},$ and $r_{C},$ the following conditions are equivalent.
@@ 3 Pythagorean and Pell-like Equations
$(x_{\sigma(1)},x_{\sigma(2)},x_{\sigma(3)},x_{4})=(t^{2}+u^{2}-v^{2}-w^{2},2(tv+uw),2(tw-uv),\pm(t^{2}+u^{2}+v^{2}+w^{2}))$ (3.1)
for some $\sigma\in\mathfrak{S}_{3}$ and $(t,u,v,w)\in\mathbb{Z}^{4}.$
@@
The well-known Euclid’s formula for Pythagorean triples is obtained by letting $t>v>u=w=0$ and $x_{4}>0$ in (3.1).
@@
If $(a_{\sigma(1)},a_{\sigma(2)},a_{\sigma(3)})=(a,0,0)$ and $(b_{\tau(1)},b_{\tau(2)},b_{\tau(3)})=(t^{2}+u^{2}-v^{2}-w^{2},2(tv+uw),2(tw-uv)),$ where $\sigma,$ $\tau\in\mathfrak{S}_{3},$ $a\in\mathbb{Z}\setminus\{0\},$ and $(t,u,v,w)\in\mathbb{Z}^{4}\setminus\{(0,0,0,0)\},$ then $\bm{a}=(a_{1},a_{2},a_{3}),$ $\bm{b}=(b_{1},b_{2},b_{3}),$ and $\bm{c}=(t^{2}+u^{2}+v^{2}+w^{2})\bm{a}\pm a\bm{b}$ satisfy (1.1).
@@ 4 Incenters of Simplices over Fields
Let $p,$ $q,$ $r\in k^{\times}.$ An equifacial tetrahedron with vertices $(p,q,r),$ (p,q,r),(p,-q,-r), (p,q,r),(-p,q,-r), and (p,q,r)(-p,-q,r) has its incenter at $(0,0,0),$ which is $k$-rational.
@@
Consider a right regular triangular pyramid $ABCD$ with $A=(p,0,0),$ $B=(0,p,0),$ $C=(0,0,p),$ and $D=(q,q,q),$ where $p,$ $q\in k,$ $p\neq 3q,$ and $p>0.$ Since $\overrightarrow{DB}=(q,q-p,q),$ $\overrightarrow{DC}=(q,q,q-p),$ and $\overrightarrow{DB}\times\overrightarrow{DC}=p(p-2q,q,q),$ we have $BCD]=(1/2)|\overrightarrow{DB}\times\overrightarrow{DC}|=(p/2)\sqrt{(p-2q)^{2}+2q^{2}}.$ Similarly, we have $CDA]=[DAB]=(p/2)\sqrt{(p-2q)^{2}+2q^{2}}.$ Furthermore, since triangle $ABC$ is a regular triangle with side length $p\sqrt{2},$ we have $ABC]=(\sqrt{3}/2)\cdot(p\sqrt{2})^{2}=p^{2}\sqrt{3}.$ Therefore, tetrahedron $ABCD$ has $k$-rational incenter $I$ if and only if
@@
$BCD]:[CDA]:[DAB]:[ABC]=\frac{pr}{2}\sqrt{3}:\frac{pr}{2}\sqrt{3}:\frac{pr}{2}\sqrt{3}:p^{2}\sqrt{3}=r:r:r:2p$
@@
$\overrightarrow{OI}=\frac{r(p,0,0)+r(0,p,0)+r(0,0,p)+2p(q,q,q)}{r+r+r+2p}=\frac{p(2q+r)}{2p+3r}(1,1,1),$
which implies $I$ is $k$-rational. For example, tetrahedron $ABCD$ with $A=(7,0,0),$ $B=(0,7,0),$ $C=(0,0,7),$ and $D=(1,1,1)$ has rational incenter $I=(35/23,35/23,35/23),$ since $(7-2\cdot 1)^{2}+2\cdot 1^{2}=3\cdot 3^{2}.$
@@ 5 Centers of Triangles over Fields
On $\mathbb{R}^{2},$ with respect to triangle $ABC$ over $k,$ the isotomic conjugate point $Q$ of a given $k$-rational point $P$ is also $k$-rational.
@@
On $\mathbb{R}^{2},$ with respect to triangle $ABC$ over $k,$ the isogonal conjugate point $Q$ of a given $k$-rational point $P$ is also $k$-rational.
@@
On $\mathbb{R}^{2},$ the centroid $G,$ the Lemoine point $L,$ the circumcenter $E,$ and the orthocenter $H$ of triangle $ABC$ over $k$ are $k$-rational.
@@
On $\mathbb{R}^{2},$ the circumradius $R$ and the inradius $r$ of Heronian triangle $ABC$ over $k$ lie in $k.$
@@
Every triangle $ABC$ over $k$ with $k$-rational incenter can be constructed by the following steps.
@@
In the following cases, the centroid $G,$ the Lemoine point $L,$ the circumcenter $E,$ the orthocenter $H,$ the incenter $I,$ the excenters $I_{A},$ $I_{B},$ and $I_{C},$ the Gergonne point $Ge,$ and the Nagel point $Na$ of triangle $ABC$ are $k$-rational. For readers’ convenience, we add the value of the vectors $\bm{u}=\overrightarrow{BC},$ $\bm{v}=\overrightarrow{CA},$ and $\bm{w}=\overrightarrow{AB},$ the side lengths $a,$ $b,$ and $c,$ the inradius $r,$ the exradii $r_{A},$ $r_{B},$ and $r_{C},$ the positive algebraic integer $d\in O_{k}$ such that $a^{2}\equiv b^{2}\equiv c^{2}\equiv d\pmod{k^{\times 2}}$, and the area $\Delta$ (see Figure 2 for the example in the middle column).

### v4

@@ Abstract
In this article, we generalize the following problem, which is called the rational angle bisection problem, to the $n$-dimensional space $k^{n}$ over a subfield $k$ of $\mathbb{R}$: in the coordinate plane, for which rational numbers $a$ and $b$ are the slopes of the angle bisectors between the two lines with slopes $a$ and $b$ rational? First, we give a few characterizations of when the angle bisectors between two lines with direction vectors in $k^{n}$ have direction vectors in $k^{n}.$ To find solutions to the problem in the case when $k=\mathbb{Q},$ we also give a formula for the integral solutions of $x_{1}{}^{2}+\dots+x_{n}{}^{2}=dx_{n+1}{}^{2}$, which is a generalization of the negative Pell’s equation $x^{2}-dy^{2}=-1,$ where $d$ is a square-free positive integer. Second, by applying the above characterizations, we give a necessary and sufficient condition for the incenter of a given $n$-simplex with $k$-rational vertices to be $k$-rational. In the coordinate plane, we prove that every triangle with $k$-rational vertices and incenter can be obtained by scaling a triangle with $k$-rational side lengths and area, which is a generalization of a Heronian triangle. We also state certain fundamental properties of a few centers of a given triangle with $k$-rational vertices.
@@ 1 Introduction
Throughout this article, let $n\geq 2$ be an integer, and let $k$ be a subfield of $\mathbb{R}.$ We consider the following problem in $\mathbb{R}^{n},$ which is a generalization of the rational angle bisection problem in $\mathbb{R}^{2}$ (see [3]).
@@
For which $\bm{a},$ $\bm{b}\in k^{n}$ do the angle bisectors between the two lines with direction vectors $\bm{a}$ and $\bm{b}$ have direction vectors in $k^{n}$?
@@
For which $\bm{a},$ $\bm{b}\in O_{k}{}^{n}$ do the angle bisectors between the two lines with direction vectors $\bm{a}$ and $\bm{b}$ have direction vectors in $O_{k}{}^{n}$?
@@
      One of the angle bisectors between the two lines with direction vectors $\bm{a}$ and $\bm{b}$ has direction vector $\bm{c}\in k^{n}.$
@@
      One of the angle bisectors between the two hyperplanes with normal vectors $\bm{a}$ and $\bm{b}$ has normal vector $\bm{c}\in k^{n}.$
@@
      One of the angle bisectors between the two lines with direction vectors $\bm{a}$ and $\bm{b}$ has direction vector $\bm{c}\in O_{k}{}^{n}$.
@@
      One of the angle bisectors between the two hyperplanes with normal vectors $\bm{a}$ and $\bm{b}$ has normal vector $\bm{c}\in O_{k}{}^{n}$.
@@
    If $\bm{a}=(2,1,1)$ and $\bm{b}=(3,6,2),$ then the angle bisectors between the two lines with direction vectors $\bm{a}$ and $\bm{b}$ have direction vectors $7\bm{a}\pm 3\bm{b},$ since $|\bm{a}|=3$ and $|\bm{b}|=7$ imply that $7\bm{a}$ and $3\bm{b}$ span a rhomb (see Figure 1).
@@
    If $\bm{a}=(1,1,4)$ and $\bm{b}=(3,4,5),$ then the angle bisectors between the two lines with direction vectors $\bm{a}$ and $\bm{b}$ have direction vectors $5\bm{a}\pm 3\bm{b},$ since $|\bm{a}|=3\sqrt{2}$ and $|\bm{b}|=5\sqrt{2}$ imply that $5\bm{a}$ and $3\bm{b}$ span a rhomb.
@@
In $\mathbb{R}^{2},$ given triangle $ABC$ over $k$ with side lengths $BC=a,$ $CA=b,$ and $AB=c,$ inradius $r,$ and exradii $r_{A},$ $r_{B},$ and $r_{C},$ the following conditions are equivalent.
@@ 3 Pythagorean and Pell-like Equations
$(x_{\sigma(1)},x_{\sigma(2)},x_{\sigma(3)},x_{4})=(s(t^{2}+u^{2}-v^{2}-w^{2}),s\cdot 2(tv+uw),s\cdot 2(tw-uv),\pm s(t^{2}+u^{2}+v^{2}+w^{2}))$ (3.1)
for some $\sigma\in\mathfrak{S}_{3}$ and $(s,t,u,v,w)\in\mathbb{Z}^{5}.$
@@
The well-known Euclid’s formula for Pythagorean triples is obtained by letting $s>0,$ $t>v>u=w=0$ and $x_{4}>0$ in (3.1).
@@
If $(a_{\sigma(1)},a_{\sigma(2)},a_{\sigma(3)})=(1,0,0)$ and $(b_{\tau(1)},b_{\tau(2)},b_{\tau(3)})=(t^{2}+u^{2}-v^{2}-w^{2},2(tv+uw),2(tw-uv)),$ where $\sigma,$ $\tau\in\mathfrak{S}_{3}$ and $(t,u,v,w)\in\mathbb{Z}^{4}\setminus\{(0,0,0,0)\},$ then $\bm{a}=(a_{1},a_{2},a_{3}),$ $\bm{b}=(b_{1},b_{2},b_{3}),$ and $\bm{c}=(t^{2}+u^{2}+v^{2}+w^{2})\bm{a}\pm\bm{b}$ satisfy (1.1).
@@ 4 Incenters of Simplices over Fields
Let $p,$ $q,$ $r\in k^{\times}.$ The equifacial tetrahedron with vertices $(p,q,r),$ (p,q,r),(p,-q,-r), (p,q,r),(-p,q,-r), and (p,q,r)(-p,-q,r) has a $k$-rational incenter at $(0,0,0).$
Let $[XYZ]$ denote the area of triangle $XYZ.$
@@
Let $p,$ $q\in k$ with $p\neq 3q$ and $p>0.$ Consider the right regular triangular pyramid $ABCD$ with $A=(p,0,0),$ $B=(0,p,0),$ $C=(0,0,p),$ and $D=(q,q,q).$ Since $\overrightarrow{DB}=(q,q-p,q),$ $\overrightarrow{DC}=(q,q,q-p),$ and $\overrightarrow{DB}\times\overrightarrow{DC}=p(p-2q,q,q),$ we have $BCD]=(1/2)|\overrightarrow{DB}\times\overrightarrow{DC}|=(p/2)\sqrt{(p-2q)^{2}+2q^{2}}.$ Similarly, we have $CDA]=[DAB]=(p/2)\sqrt{(p-2q)^{2}+2q^{2}}.$ Furthermore, since triangle $ABC$ is the regular triangle with side length $p\sqrt{2},$ we have $ABC]=(\sqrt{3}/4)\cdot(p\sqrt{2})^{2}=p^{2}\sqrt{3}/2.$ Therefore, tetrahedron $ABCD$ has a $k$-rational incenter $I$ if and only if
@@
$BCD]:[CDA]:[DAB]:[ABC]=\frac{pr}{2}\sqrt{3}:\frac{pr}{2}\sqrt{3}:\frac{pr}{2}\sqrt{3}:\frac{p^{2}}{2}\sqrt{3}=r:r:r:p$
@@
$\overrightarrow{OI}=\frac{r(p,0,0)+r(0,p,0)+r(0,0,p)+p(q,q,q)}{r+r+r+p}=\frac{p(q+r)}{p+3r}(1,1,1),$
which implies $I$ is $k$-rational. For example, tetrahedron $ABCD$ with $A=(7,0,0),$ $B=(0,7,0),$ $C=(0,0,7),$ and $D=(1,1,1)$ has a rational incenter $I=(7/4,7/4,7/4),$ since $(7-2\cdot 1)^{2}+2\cdot 1^{2}=3\cdot 3^{2}.$
@@ 5 Centers of Triangles over Fields
In $\mathbb{R}^{2},$ with respect to triangle $ABC$ over $k,$ the isotomic conjugate point $Q$ of a given $k$-rational point $P$ is also $k$-rational.
@@
In $\mathbb{R}^{2},$ with respect to triangle $ABC$ over $k,$ the isogonal conjugate point $Q$ of a given $k$-rational point $P$ is also $k$-rational.
@@
In $\mathbb{R}^{2},$ the centroid $G,$ the Lemoine point $L,$ the circumcenter $E,$ and the orthocenter $H$ of triangle $ABC$ over $k$ are $k$-rational.
@@
In $\mathbb{R}^{2},$ the circumradius $R$ and the inradius $r$ of Heronian triangle $ABC$ over $k$ lie in $k.$
@@
In $\mathbb{R}^{2},$ every triangle $ABC$ over $k$ with $k$-rational incenter can be constructed by the following steps.
@@
In the following cases, the centroid $G,$ the Lemoine point $L,$ the circumcenter $E,$ the orthocenter $H,$ the incenter $I,$ the excenters $I_{A},$ $I_{B},$ and $I_{C},$ the Gergonne point $Ge,$ and the Nagel point $Na$ of triangle $ABC$ are $k$-rational. For readers’ convenience, we add the values of the vectors $\bm{u}=\overrightarrow{BC},$ $\bm{v}=\overrightarrow{CA},$ and $\bm{w}=\overrightarrow{AB},$ the side lengths $a,$ $b,$ and $c,$ the inradius $r,$ the exradii $r_{A},$ $r_{B},$ and $r_{C},$ the positive algebraic integer $d\in O_{k}$ such that $a^{2}\equiv b^{2}\equiv c^{2}\equiv d\pmod{k^{\times 2}}$, and the area $\Delta$ (see Figure 2 for the example in the middle column).